A population consists of the numbers 2, 4, 5, 9, and 10. A random sample of size 2 is


taken from the population.


1. Compute the number of samples using combination.


2. Complete the table on the right and compute the mean µ, variance


σ2 and standard deviation σ of the population.


3. Complete the table on the right and compute


the mean μx̄, variance σx̄2


, and standard


deviation σx̄ of the sampling distribution of


sample means.

1
Expert's answer
2022-05-16T15:27:29-0400

We have population values 2,4,5,9,10, population size N=5 and sample size n=2.

Mean of population "(\mu)" = "\dfrac{2+4+5+9+10}{5}=6"

Variance of population 


"\sigma^2=\dfrac{\Sigma(x_i-\bar{x})^2}{n}=\dfrac{16+4+1+9+16}{5}=9.2"


"\sigma=\sqrt{\sigma^2}=\sqrt{9.2}\approx3.03315"

A. Select a random sample of size 2 without replacement. We have a sample distribution of sample mean.

The number of possible samples which can be drawn without replacement is "^{N}C_n=^{5}C_2=10."

"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 no & Sample & Sample \\
& & mean\ (\bar{x})
\\ \hline
 1 & 2,4 & 6/2 \\
 \hdashline
 2 & 2,5 & 7/2 \\
 \hdashline
 3 & 2,9 & 11/2 \\
 \hdashline
 4 & 2,10 & 12/2 \\
 \hdashline
 5 & 4,5 & 9/2 \\
 \hdashline
 6 & 4,9 & 13/2 \\
 \hdashline
 7 & 4,10 & 14/2 \\
 \hdashline
 8 & 5,9 & 14/2 \\
 \hdashline
 9 & 5,10 & 15/2 \\
 \hdashline
 10 & 9,10 & 19/2 \\
 \hdashline
\end{array}"





"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 \bar{X} & f(\bar{X}) &\bar{X} f(\bar{X}) & \bar{X}^2f(\bar{X})
\\ \hline
 6/2 & 1/10 & 6/20 & 36/40 \\
 \hdashline
 7/2 & 1/10 & 7/20 & 49/40 \\
 \hdashline
 9/2 & 1/10 & 9/20 & 81/40 \\
 \hdashline
 11/2 & 1/10 & 11/20 & 121/40 \\
 \hdashline
 12/2 & 1/10 & 12/20 & 144/40 \\
 \hdashline
 13/2 & 1/10 & 13/20 & 169/40 \\
 \hdashline
 14/2 & 2/10 & 28/20 & 392/40 \\
 \hdashline
 15/2 & 1/10 & 15/20 & 225/40 \\
 \hdashline
 19/2 & 1/10 & 19/20 & 361/40 \\
 \hdashline
\end{array}"



Mean of sampling distribution 

"\mu_{\bar{X}}=E(\bar{X})=\sum\bar{X}_if(\bar{X}_i)=\dfrac{120}{20}=6=\mu"



The variance of sampling distribution 

"Var(\bar{X})=\sigma^2_{\bar{X}}=\sum\bar{X}_i^2f(\bar{X}_i)-\big[\sum\bar{X}_if(\bar{X}_i)\big]^2""=\dfrac{1578}{40}-(6)^2=3.45= \dfrac{\sigma^2}{n}(\dfrac{N-n}{N-1})"

"\sigma_{\bar{X}}=\sqrt{3.45}\approx1.8574"